# MaRI: Material Retrieval Integration across Domains

### v1

@@ 1 Introduction
To handle these challenges, we introduce MaRI—a novel framework inspired by CLIP that learns a joint embedding space for both materials and images. MaRI employs dual encoders, jointly trained to align material properties with visual features in a shared space, facilitating direct and efficient comparisons. By leveraging pre-trained DINOv2 models as the backbone for both the material and image encoders, MaRI preserves generalizability while fine-tuning only the final Transformer block to capture domain-specific nuances effectively. We construct a large-scale dataset pairing images with materials, designed to capture both diversity and realism, for training a joint embedding. Since such a dataset is unavailable, we adopt both synthesis and generative approach to automatically construct the dataset. During synthesis, we construct each data pair by sampling a material from a material gallery, associating it to an object from an object dataset, and render it with Blender to obtain an image that pairs with the material. Although synthetic data alone yields satisfactory results for training, it still partially falls short in bridging the domain gap, as it cannot fully represent the diverse and nuanced appearances of real-world materials. To complement this, we introduce a generative approach that incorporate large-scale, unlabeled real-world image data and construct paired material with a material transfer technique (ZeST) [7]. As a result, the generative approach captures diverse material appearances under varying conditions through real images. It enables us to adopt an unsupervised learning approach, helping the model to learn robust material representations without being dependent on annotated datasets.
@@ 2 Related Work
Material Generation and Retrieval. Advances in material generation and retrieval have enhanced the realism of 3D content creation. Techniques like ControlMat [33] use diffusion models to generate high-resolution material maps, allowing precise control over surface properties such as roughness and normal maps. MatFuse [36] enables users to guide the material generation process through sketches, color palettes, or text prompts, providing greater flexibility in design. Similarly, Fantasia3D [6] employs pixel-level optimization for detailed material representations, though it faces challenges in maintaining stability with complex geometries. Make-it-3D [30] uses a two-stage diffusion process to generate high-fidelity 3D content from single images, emphasizing texture refinement for realistic outputs. Tools like MaterialPlatte [23] and Matlas [4] further advance material manipulation by offering more control over varied environmental conditions and geometries.
MaPa [39] and Make-it-Real [14] integrate material retrieval within broader 3D generation workflows. MaPa utilizes GPT-4V [13] for initial material categorization and CLIP [26] for refining the material assignment process, retrieving material graphs from predefined libraries, though it often lacks precision in handling fine textures and complex surfaces. Make-it-Real leverages GPT-4V to assign materials to segmented 3D models using SVBRDF [16] mappings, but its reliance on pre-annotated datasets limits adaptability to novel materials and environments. These limitations underscore the need for more robust frameworks. Our work addresses this by bridging the gap between real-world and synthetic data, capturing both fine textures and complex material details.
@@ 3.3 Domain-Adaptive Contrastive Learning
Building on the diverse range of material data in $\mathcal{D}$, our proposed MaRI framework is inspired by the contrastive learning approach of CLIP, but adapts it to bridge the domain gap between synthetic and real-world visual representations—a key challenge in material retrieval tasks. Rather than aligning different modalities, MaRI focuses on aligning varied visual features within a single modality but across different domains. It utilizes two DINOv2-based encoders, $E_{r}$ for masked image representations and $E_{m}$ for material properties, to project masked image inputs $x_{i}$ and material spheres $m_{i}$ into a shared feature space $\mathcal{F}$:
$\mathbf{z}_{r}^{i}=E_{r}(x_{i}\odot\text{mask}_{i}),\quad\mathbf{z}_{m}^{i}=E_{m}(m_{i}).$ (4)
Here, $\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{i}\in\mathbb{R}^{d}$ are the embeddings of the masked rendered image and the material sphere image, and $\odot$ denotes element-wise multiplication to apply the mask.
@@
$s(\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{j})=\frac{\mathbf{z}_{r}^{i}\cdot\mathbf{z}_{m}^{j}}{\sqrt{d}},$ (5)
@@
$\mathcal{L}_{\text{contrast}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\bigl{(}s(\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{i})/\tau\bigr{)}}{\sum_{j=1}^{N}\exp\bigl{(}s(\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{j})/\tau\bigr{)}}.$ (6)
The loss encourages positive pairs $(\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{i})$ to have higher similarity scores than any other pair $(\mathbf{z}_{r}^{i},\mathbf{z}_{m}^{j})$ with $j\neq i$, thus aligning the features in a domain-agnostic manner. MaRI effectively creates a shared space that supports robust material retrieval across varying data sources.
@@ 4.2 Comparative Analysis of Material Retrieval
The qualitative results in Figure 4 illustrate that general image search models, such as the original DINOv2, struggle to capture the intricate relationships between material textures and their corresponding images. As previously demonstrated through quantitative evaluations, methods like CLIP and ViT exhibit similarly poor performance and are therefore omitted from this figure. By contrast, MaRI outperforms GPT-4V-based material retrieval methods, including MaPa and Make-it-Real, by more effectively capturing fine-grained material characteristics. For instance, in the ”Bark” case shown in Figure 4, MaRI consistently retrieves materials in its top-5 predictions that closely resemble the target in both texture and color.

### v3

# SuppSupplementary References

@@ 1 Introduction
To handle these challenges, we introduce MaRI—a novel framework inspired by CLIP that learns a joint embedding space for both materials and images. MaRI employs dual encoders, jointly trained to align material properties with visual features in a shared space, facilitating direct and efficient comparisons. By leveraging pre-trained DINOv2 models as the backbone for both the material and image encoders, MaRI preserves generalizability while fine-tuning only the final Transformer block to capture domain-specific nuances effectively. We construct a large-scale dataset pairing images with materials, designed to capture both diversity and realism, for training a joint embedding. Since such a dataset is unavailable, we adopt both synthesis and generative approach to automatically construct the dataset. During synthesis, we construct each data pair by sampling a material from a material gallery, associating it to an object from an object dataset, and render it with Blender to obtain an image that pairs with the material. Although synthetic data alone yields satisfactory results for training, it still partially falls short in bridging the domain gap, as it cannot fully represent the diverse and nuanced appearances of real-world materials. To complement this, we introduce a generative approach that incorporate large-scale, unlabeled real-world image data and construct paired material with a material transfer technique (ZeST) [7]. As a result, the generative approach captures diverse material appearances under varying conditions through real images. It enables us to adopt a self-supervised learning framework, helping the model to learn robust material representations without being dependent on annotated datasets.
@@ 2 Related Work
Material Generation and Retrieval. Advances in material generation and retrieval have enhanced the realism of 3D content creation. Techniques like ControlMat [41] use diffusion models to generate high-resolution material maps, allowing precise control over surface properties such as roughness and normal maps. MatFuse [42] enables users to guide the material generation process through sketches, color palettes, or text prompts, providing greater flexibility in design. Similarly, Fantasia3D [6] employs pixel-level optimization for detailed material representations, though it faces challenges in maintaining stability with complex geometries. Make-it-3D [37] uses a two-stage diffusion process to generate high-fidelity 3D content from single images, emphasizing texture refinement for realistic outputs. Tools like Material Palette [28] and Matlas [4] further advance material manipulation by offering more control over varied environmental conditions and geometries.
Several studies [22, 34] have focused on perceptual similarity. However, none directly address the challenge of retrieving accurate materials from real-world images. MaPa [45] and Make-it-Real [14] integrate material retrieval within broader 3D generation workflows. MaPa utilizes GPT-4V [13] for initial material categorization and CLIP [31] for refining the material assignment process, retrieving material graphs from predefined libraries, though it often lacks precision in handling fine textures and complex surfaces. Make-it-Real leverages GPT-4V to assign materials to segmented 3D models using SVBRDF [16] mappings, but its reliance on pre-annotated datasets limits adaptability to novel materials and environments. These limitations underscore the need for more robust frameworks. Our work addresses this by bridging the gap between real-world and synthetic data, capturing both fine textures and complex material details.
@@ 3.3 Domain-Adaptive Contrastive Learning
Building on the diverse range of material data in $\mathcal{D}$, our proposed MaRI framework is inspired by the contrastive learning approach of CLIP, but adapts it to bridge the domain gap between synthetic and real-world visual representations—a key challenge in material retrieval tasks. Rather than aligning different modalities, MaRI focuses on aligning varied visual features within a single modality but across different domains. It utilizes two DINOv2-based encoders, $E_{I}$ for masked image representations and $E_{M}$ for material properties, to project masked image inputs $x_{i}$ and material spheres $m_{i}$ into a shared feature space $\mathcal{F}$:
$\mathbf{z}_{I}^{i}=E_{I}(x_{i}\odot\text{mask}_{i}),\quad\mathbf{z}_{M}^{i}=E_{M}(m_{i}).$ (4)
Here, $\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{i}\in\mathbb{R}^{d}$ are the embeddings of the masked rendered image and the material sphere image, and $\odot$ denotes element-wise multiplication to apply the mask.
@@
$\text{sim}(\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{j})=\frac{\mathbf{z}_{I}^{i}\cdot\mathbf{z}_{M}^{j}}{\sqrt{d}},$ (5)
@@
$\mathcal{L}_{\text{contrast}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\bigl{(}\text{sim}(\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{i})/\tau\bigr{)}}{\sum_{j=1}^{N}\exp\bigl{(}\text{sim}(\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{j})/\tau\bigr{)}}.$ (6)
The loss encourages positive pairs $(\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{i})$ to have higher similarity scores than any other pair $(\mathbf{z}_{I}^{i},\mathbf{z}_{M}^{j})$ with $j\neq i$, thus aligning the features in a domain-agnostic manner. MaRI effectively creates a shared space that supports robust material retrieval across varying data sources.
@@ 4.2 Comparative Analysis of Material Retrieval
The qualitative results in Figure 4 illustrate that general image search models, such as the original DINOv2, struggle to capture the intricate relationships between material textures and their corresponding images. As previously demonstrated through quantitative evaluations, methods like CLIP and ViT exhibit similarly poor performance and are therefore omitted from this figure. In contrast, MaRI also outperforms GPT-4V-based material retrieval methods, including MaPa and Make-it-Real, by more effectively capturing fine-grained material characteristics. For instance, in the “Bark” case shown in Figure 4, MaRI consistently retrieves materials in its Top-5 predictions that closely resemble the target in both texture and color.
@@ 5 Conclusion
Supplementary Material
This supplementary material provides additional details and insights into the MaRI framework. Section 1 outlines the implementation details, including training configurations and the use of pre-trained DINOv2 [29] backbones. Section 2 presents an analysis of how different backbone architectures impact material retrieval performance on both trained and unseen datasets. Practical applications of MaRI, such as assigning materials to 3D models for design workflows, are demonstrated in Section 3, showcasing the system’s versatility and ease of use. In addition, Section 4 reports a user study, and Section 5 offers further discussion on limitations, and potential future improvements.
1 Implementation Details
We use DINOv2 as the backbone for both the image and material encoders, initialized with pre-trained weights. The training process consists of two main stages: first, the model is fine-tuned on the synthetic dataset for 1 epoch using the Adam optimizer with a learning rate of $1\times 10^{-4}$. This initial phase helps establish a robust baseline. Next, the model undergoes fine-tuning on the real-world dataset for 25 epochs at a reduced learning rate of $1\times 10^{-5}$ to enhance generalization by capturing more intricate real-world features. To facilitate effective alignment in the shared feature space, the temperature parameter $\tau$ in the contrastive loss is set to 0.07. The dataset is divided into 90% for training and 10% for validation, and a batch size of 256 is employed. The entire process is conducted on four NVIDIA A100 GPUs (80GB each), completing within 3 hours.
2 Effect of Backbone Variations
The performance of different backbone architectures for material retrieval is summarized in Table 1. DINOv2 demonstrates the strongest overall performance, achieving top-1 instance accuracy (T1I) of 26.0% and top-5 instance accuracy (T5I) of 90.0% on the Trained dataset, along with 54.0% and 89.0% on the Unseen dataset, respectively.
Table 1: Backbone comparison for material retrieval on Trained and Unseen datasets. Best values are highlighted in blue.
Backbone Trained Unseen
T1I T5I T1C T3IoU T1I T5I
ResNet50 [19] 7.5% 28.0% 64.5% 0.58 30.0% 60.5%
ViT [12] 15.0% 33.5% 54.5% 0.67 21.0% 65.0%
Swin Transformer [26] 14.0% 41.5% 77.0% 0.68 38.5% 77.0%
ConvNeXt [27] 16.5% 42.0% 70.0% 0.67 35.0% 73.5%
EfficientNet [36] 8.5% 21.5% 52.0% 0.55 22.0% 54.0%
CLIP [31] 12.0% 44.5% 61.0% 0.68 29.5% 72.5%
DINOv2 [29] 26.0% 90.0% 81.5% 0.77 54.0% 89.0%
Table 1 highlights DINOv2’s capability to effectively capture complex material characteristics and generalize across diverse data distributions. Other architectures show moderate performance, reflecting limited generalization capacity. These findings emphasize the critical role of backbone selection in enhancing material retrieval tasks, with DINOv2 emerging as the most effective backbone for bridging the synthetic-to-real domain gap.
3 Applications
Having established the effectiveness of MaRI in retrieving materials accurately, we now highlight a practical application that leverage MaRI’s capabilities in real-world 3D design workflows. MaRI empowers users to effortlessly assign desired materials to different parts of a 3D model by providing a streamlined paradigm for material retrieval and application. For instance, a user working with a 3D chair model can specify preferred materials—such as leather for the seat, metal for the legs, and wood for the armrests—through simple input reference images. MaRI efficiently retrieves matching physically based rendering (PBR) materials and applies them to the respective parts. The workflow involves the following steps:
  1. 1.
    Object Segmentation: The 3D model is segmented into distinct components (e.g., seat, legs, and armrests) based on its structural design or user annotations.
  2. 2.
    Material Retrieval: Users provide reference images for their desired material patterns, such as leather textures or wood grains. MaRI leverages its robust library to retrieve accurate and visually consistent matches.
  3. 3.
    Material Application: The retrieved materials are mapped and applied to the segmented parts of the model, enabling users to visualize their design with photorealistic fidelity.
This process, exemplified through the chair model in Figure 2, highlights MaRI’s ability to offer users an intuitive and efficient workflow for realizing their creative goals. MaRI transforms the traditionally labor-intensive process, allowing users to quickly locate and apply desired materials, resulting in outcomes that closely match their creative intent.
Figure 1: Comparison between Material Palette and MaRI.
4 User Study
We conducted a user study with 20 participants (40% designers, 30% researchers, and 30% general users). Each participant provided three material images as queries, and the top-5 retrievals from MaRI and baseline methods were rated on relevance, realism, and perceptual consistency using a 5-point Likert scale. Table 2 shows that MaRI achieved consistently higher scores across all top-k retrieval results. These results indicate that our method retrieves materials that are more relevant, realistic, and perceptually consistent compared to the baselines.
Table 2: User study average scores for top-k retrievals. Best values are highlighted in blue.
Method Top-1 Top-3 Top-5
MaPa [45] 2.60 3.05 3.80
Make-it-Real [14] 3.25 3.50 3.90
MaRI 4.15 4.50 4.85
5 Discussion
We further analyze the behavior of our retrieval system. As shown in Figure 7, the last row presents a case where a query for sponge material returns several sand-like materials. This outcome is due to the limited number of sponge samples available in our gallery, which causes the system to favor visually similar textures. The case highlights a potential limitation of our approach compared to direct generative methods like Material Palette [28], which can synthesize a broader variety of material appearances. However, our retrieval-based approach is faster than diffusion-based methods. While Material Palette generates novel material images using diffusion, our method retrieves assets from an existing database. This distinction is crucial, as it ensures that our results are not only of high quality but also readily support PBR rendering. Figure 1 shows a comparison case.
In our work, we adopt material spheres as the shape representation, which is a widely used choice among content creators. MaRI effectively captures essential material properties and supports robust retrieval performance. We believe that exploring alternative representations, such as blob [39] or Havran’s [18] shapes, may further enhance perceptual alignment [34] by capturing subtle material nuances. This presents a promising direction for future research.
Figure 2: Material assignment and rendering for a 3D chair model.
Figure 3: Material assignment and rendering for a 3D plant model.
Figure 4: Material assignment and rendering for a 3D desk model.
Figure 5: Material assignment and rendering for a 3D bucket model.
Figure 6: Top-5 material retrieval results for real-world images.
Figure 7: Top-5 material retrieval results for real-world images.